Page 1 of 1
linear velocity problem
Posted: Mon Jan 15, 2007 12:15 pm
by rajdeeproy
Dear Peter , bruce and all
My method flow rate is 5 ml/ min and the pressure used to be 8.4 psi. The linear velocity of varied between 32- 34 cms/sec. The flow rate is EFC controlled and i cant controlled this pressure it is automatically shown.
I am using the Varian system CP 3800. Recently i had UPS shut down for routine maintinence after that the pressure is changed to 84 PSI and liner velocity is fluctuating, but the flow rate is constant.
The retention time has also changed.
please give me some advice
rajdeep
Posted: Mon Jan 15, 2007 2:58 pm
by Peter Apps
Hi Rajdeep
I assume from the linear and volume flows that you used to have that you are using a 530 micron column. If not you have a very low split ratio into a 320 or 250 micron column.
My guess is that when the power was off the GC lost its initial settings - check in the instruction book how to initialise the flow sensors on start up.
If your volume flow (ml/min) is constant, the linear volcity cannot fluctuate.
84 psi is a very high pressure for a GC - are you sure that it is not KPa ?
Peter
Posted: Mon Jan 22, 2007 7:42 am
by rajdeeproy
Dear peter and all
I tried to do EFC calibration and again the pressure increased. Then i removed the column and changed the ferule and things are fine. I strongly feel there is some problem with the EFC.The base line has also increased. It used to be at 10mv, now it has increased to 23mv.
I should mention that i dont do autozero during my analysis. The current to the ecd is adjusted to very low, known as NTSTD in varian GCs. According to the Varian manual a base line above 25 mv decreases the linear response of the detector. This all happens when i keep at range 1. In range 10 the base line is 1.0 mv.
please do send me your suggestion
Posted: Tue Jan 23, 2007 6:47 am
by Peter Apps
We need full details of your set up; column dimensions, type of inlet, split ratio if any etc etc etc etc
full details of the method
Posted: Thu Jan 25, 2007 11:11 am
by rajdeeproy
Dear peter and all
here is the full details.
column (DB624+HP5 0.53mmID)) 60m+15m connected by fused silica joints (supelco)
column flow 5ml/min (He)
oven temperature programing
40 degree for one min
raised to 180 degree@ 4degree/ min
hold 25 min
injector 110
detector ECD 230 degree
make up nitrogen 30 ml / min
if you want more details please let me know
Posted: Thu Jan 25, 2007 12:38 pm
by Peter Apps
Hi Rajdeep
On the face of it there is nothing strange about your set up.
What is the pressure reading now ?, if you inject some headspace of dichloromethane how long does it take the peak to appear ?.
If things are fine after you changed the ferrule, why do you think that there is something wrong with the EFC ?.
Have you checked for leaks - with a leak seeker ?
Which EFC module do you have (1 - 4), is it set for flow or pressure ?
Peter
Posted: Thu Jan 25, 2007 1:13 pm
by rajdeeproy
Dear peter,
two days back it was showing 8.2 psi and now it again started fluctuating now it is showing 61psi but flow rate same as 5ml/min.I will check by injecting headspace of ch2cl2.
The type of EFC is 3 , you can change flow rate but not pressure. U need to give details of carrier gas and column and linear velocity is automatically calculated and corrosponding pressure is shown.
I do run a hplc simultaneously i dont get any problem. GC is just becoming a big deal for me.
I will send u more details of ECD cell current etc.
I do have another question. Do u do autozero while running ur samples?
best regards
rajdeep
Posted: Thu Jan 25, 2007 2:04 pm
by Peter Apps
Hi Rajdeep
It is entirely possible that the EFC is malfunctioning - after you have tested the retention of dichloromethane you could remove the column from the detector (plug the detector connection) and measure the volume flow through the column with a bubble flow meter.
You can also double-check the inlet pressure by connecting a syringe needle to a pressure gauge and sticking it through the septum.
If the flow and/or pressure really is fluctuating as much as it seems to be it would be enough to cause problems with the detector baseline.
Peter
Posted: Thu Feb 01, 2007 1:38 pm
by rajdeeproy
Dear peter,
There was a problem with the fused silica connector. After removing the head pressure is now 8.2 psi.